### v1

@@ 3.2 Aubin property implies SSOSC
Since the feasible set of (3.21) is compact, the minimum of the objective function can be attained at a certain solution $(\bar{\eta};\bar{\Theta}_{1};\dots;\bar{\Theta}_{J})$. Moreover, it is easy to see that the optimal value of (3.21) is not positive since (3.20) does not hold. By applying Lemma 2.4 with ${\mathcal{H}}={\mathcal{A}},\,K=-({\mathcal{S}}^{|\beta_{1}|}_{+}\times\dots\times{\mathcal{S}}^{|\beta_{J}|}_{+}),\,{\mathcal{E}}={\mathds{R}}^{r}$ and ${\mathcal{F}}={\mathcal{S}}^{|\beta_{1}|}\times\dots\times{\mathcal{S}}^{|\beta_{J}|}$, one has that Robinson’s constraint qualification holds for the constraint in (3.21) at $(\bar{\eta};\bar{\Theta}_{1};\dots;\bar{\Theta}_{J})$. Consequently, we know from [6, Theorem 3.9] that there exists a Lagrange multiplier $(\bar{\rho};\bar{\tau};\bar{\Delta}_{1};\dots;\bar{\Delta}_{J})\in{\mathds{R}}^{r}\times{\mathds{R}}\times{\mathcal{S}}^{|\beta_{1}|}\times\dots\times{\mathcal{S}}^{|\beta_{J}|}$ at $(\bar{\eta};\bar{\Theta}_{1};\dots;\bar{\Theta}_{J})$ such that the following KKT system of (3.21) holds:
$\begin{cases}({\mathcal{W}}^{*}{\mathcal{Q}}{\mathcal{W}})^{-1}\bar{\eta}+\bar{\rho}=0,\\[2.84526pt] -(P_{j})_{\beta_{j}}^{\top}(g_{j}^{\prime}(\bar{x}){\mathcal{W}}\bar{\rho})(P_{j})_{\beta_{j}}+\bar{\tau}\bar{\Theta}_{j}+\bar{\Delta}_{j}=0,\\[2.84526pt] \bar{\eta}-\sum_{j=1}^{J}{\mathcal{W}}^{*}\nabla g_{j}(\bar{x})(P_{j})_{\beta_{j}}\bar{\Theta}_{j}(P_{j})_{\beta_{j}}^{\top}=0,\\[2.84526pt] \frac{1}{2}(\sum_{j=1}^{J}\|\bar{\Theta}_{j}\|^{2}-1)=0,\\[2.84526pt] \bar{\Theta}_{j}\preceq 0,\,\bar{\Delta}_{j}\succeq 0,\langle\bar{\Delta}_{j},\bar{\Theta}_{j}\rangle=0,\;j=1,\dots,J.\end{cases}$ (3.22)
Note that
@@
$\bar{\tau}$ $=\langle(\bar{\Delta}_{1};\dots;\bar{\Delta}_{J})+\bar{\tau}(\bar{\Theta}_{1};\dots;\bar{\Theta}_{J}),(\bar{\Theta}_{1};\dots;\bar{\Theta}_{J})\rangle$
@@
$=\left\langle\bar{\rho},\sum_{j=1}^{J}{\mathcal{W}}^{*}\nabla g_{j}(\bar{x})(P_{j})_{\beta_{j}}\bar{\Theta}_{j}(P_{j})_{\beta_{j}}^{\top}\right\rangle=\langle\bar{\rho},\bar{\eta}\rangle=-\langle\bar{\eta},({\mathcal{W}}^{*}{\mathcal{Q}}{\mathcal{W}})^{-1}\bar{\eta}\rangle\geq 0,$
where the last inequality holds because the optimal value of (3.21) is not positive. Note that one has the eigenvalue decomposition that
@@
where $\bar{O}_{j}$ is an orthogonal matrix. Moreover, from the last line of (3.22) one has that
@@
$\bar{\Delta}_{j}+\bar{\tau}\bar{\Theta}_{j}=\bar{O}_{j}\begin{pmatrix}(\bar{\Lambda}_{j})_{\bar{\alpha}_{j}\bar{\alpha}_{j}}&0&0\\ 0&0_{\bar{\beta}_{j}\bar{\beta}_{j}}&0\\ 0&0&\bar{\tau}(\bar{\Lambda}_{j})_{\bar{\gamma}_{j}\bar{\gamma}_{j}}\end{pmatrix}\bar{O}_{j}^{\top},\quad j=1,\dots,J.$
For arbitrarily given $j\in\{1,\ldots,J\}$, one can take $\beta_{+}=\bar{\alpha}_{j}$, $\beta_{0}=\bar{\beta}_{j}\cup\bar{\gamma}_{j}$, and $\beta_{-}=\emptyset$ in (2.6) and (2.7) to obtain
@@
$\begin{cases}\Xi_{1}\circ\bar{O}^{\top}\bar{\Theta}_{j}\bar{O}=0=\Xi_{2}\circ\bar{O}^{\top}(\bar{\Delta}_{j}+\bar{\tau}\bar{\Theta}_{j})\bar{O},\\ \bar{O}_{\beta_{0}}^{\top}\bar{\Theta}_{j}\bar{O}_{\beta_{0}}\preceq 0,\\ \bar{O}_{\beta_{0}}^{\top}(\bar{\Delta}_{j}+\bar{\tau}\bar{\Theta}_{j})\bar{O}_{\beta_{0}}\preceq 0,\end{cases}$
which, together with Lemma 2.1 and (3.22), implies that
$\bar{\Theta}_{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{|\beta_{j}|}}(0,0)(\bar{\Delta}_{j}+\bar{\tau}\bar{\Theta}_{j})={\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{|\beta_{j}|}}(0,0)((P_{j})_{\beta_{j}}^{\top}(g_{j}^{\prime}(\bar{x}){\mathcal{W}}\bar{\rho})(P_{j})_{\beta_{j}}).$
Note that such an inclusion holds simultaneously for all $j=1,\dots,J$. Thus, this inclusion, together with (3.23), makes a contradiction to (3.12) in Proposition 3.2 (with $\nu=\bar{\rho}$ and $\widetilde{U}^{j}_{\beta_{j}\beta_{j}}=\bar{\Theta}_{j}$ for all $j=1,\ldots,J$). As a consequence, we know that (3.20) is valid. Hence, one has (3.19) holds, which leads to the SSOSC (2.18) in Definition 2.1. This completes the proof. ∎
4 Equivalent characterizations of the Aubin property
This section establishes the equivalent characterizations of the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ (1.6) at $(0;0)$ for $(\bar{x};\bar{y})$ with $\bar{x}$ being a locally optimal solution to the NLSDP problem (1.8) and $\bar{y}\in{\mathcal{M}}(\bar{x})$.
We first review some related concepts in variational analysis. As was mentioned in Section 1, the Aubin property is related to the strong metric regularity [16, Definition 2.5]. Specifically, for a set-valued mapping $\Psi:{\mathcal{E}}\rightrightarrows{\mathcal{F}}$, one has $\Psi$ is strongly metrically regular at $(\bar{z};\bar{w})\in{\sf gph}\Psi$ if $\Psi^{-1}$ has the Aubin property at $\bar{w}$ for $\bar{z}$, and there exist neighborhoods ${\mathcal{U}}$ of $\bar{z}$ and ${\mathcal{V}}$ of $\bar{w}$, such that $\Psi^{-1}(w)\cap{\mathcal{U}}$ is a singleton for all $w\in{\mathcal{V}}$. The following result provides a criterion for characterizing the strong metric regularity.
@@ 4 Equivalent characterizations of the Aubin property
[16, Theorem 2.7] $\Psi$ is strongly metrically regular at $(\bar{z};\bar{w})\in{\sf gph}(\Psi)$ if and only if for all $w\in{\mathcal{F}}$ and $z\in{\mathcal{E}}$, one has
@@
Since $F$ is locally Lipschitz continuous around $(\bar{x};\bar{y})$, it is almost everywhere differentiable in a neighborhood ${\mathcal{V}}$ of $(\bar{x};\bar{y})$ by Rademacher’s theorem [40, Theorem 9.60]. We use $D_{F}\subseteq{\mathcal{V}}$ to denote the set of points at which $F$ is differentiable. The Bouligand subdifferential of $F$ at $(\bar{x};\bar{y})$ is defined by
@@
The Clarke generalized jacobian of $F$ at $(\bar{x};\bar{y})$ is defined by
@@
$\min_{x\in{\mathcal{X}},a\in{\mathcal{X}},b\in Y}\ \phi(x,b)-\langle a,x\rangle,$ (4.2)
@@
$\begin{cases}\displaystyle M_{\iota}(a,b):=\operatorname*{arg\,min}_{x\in{\mathcal{X}}}\{\phi(x,b)-\langle a,x\rangle\mid\|x-\bar{x}\|\leq\iota\},\\ \displaystyle m_{\iota}(a,b):=\inf_{x\in{\mathcal{X}}}\{\phi(x,b)-\langle a,x\rangle\mid\|x-\bar{x}\|\leq\iota\}.\end{cases}$
We say the point $\bar{x}$ is a Lipschitzian fully stable local minimizer [28, Definition 3.2] of (4.2) at $(\bar{a};\bar{b})$ if there exist numbers $\kappa>0,\iota>0$ and a neighborhood ${\mathcal{V}}$ of $(\bar{a};\bar{b})$ such that the mapping $M_{\iota}(a,b)$ is single-valued on ${\mathcal{V}}$ with $M_{\iota}(\bar{a},\bar{b})=\bar{x}$ satisfying
@@
and that the function $m_{\iota}(a,b)$ is also Lipschitz continuous on ${\mathcal{V}}$. Different from ${\mathbb{S}}_{\rm KKT}$, the mapping $M_{\iota}(a,b)$ focuses mainly on locally optimal solutions.
Recently, Rockafellar [38, 39] introduced the strong variational sufficient condition for local optimality, and provided several characterizations of this abstract property. With $\phi$ being defined in (4.3), one can reformulate (1.1) to
@@
Define the function $\phi_{\ell}(x,b):=\phi(x,b)+\frac{\ell}{2}\|b\|^{2}$. The following definition comes from [39, Section 2].
@@
and $\varrho(x,b)=\phi_{\ell}(x,b)$ holds for $((x;b);(a;y))$ belonging to this common set. Here $\partial\varrho$ is the subdifferential in convex analysis and $\partial\phi_{\ell}$ is the limiting subdifferential defined in (2.1).
@@
For a special case ($J=1$ and $m=0$) of the NLSDP problem (1.8), the strong variational sufficient condition (with respect to a solution $(\bar{x};\bar{y})$ to the KKT system (1.3)) and the SSOSC (2.18) (for the same $(\bar{x};\bar{y})$) were proved to be equivalent in [43, Theorem 3.3]. This equivalence can be extended to the NLSDP problem (1.8) in its general form with ease by directly following their proof.
Based on the above definitions, we can establish the following result regarding the equivalent characterizations of the Aubin property for the solution mapping ${{\mathbb{S}}}_{\rm KKT}$ in (1.6) of the NLSDP problem (1.8).
@@
    The strong second-order sufficient (2.18) holds at $(\bar{x};\bar{y})$ and $\bar{x}$ is nondegenerate, i.e., (2.14) holds.
@@
One has $\rm(i)\Rightarrow\rm(ii)$ from Proposition 3.3 and [23, Theorem 1]. By simply repeating the proof of [42, Proposition 3.2] one can get $\rm(ii)\Rightarrow\rm(iii)\Rightarrow\rm(iv)$. It follows by [34, Corollary 2.2] and the definition that $\rm(iv)\Rightarrow\rm(v)$ holds. From Lemma 4.1 and the Mordukhovich criterion for the Aubin property [40, Theorem 9.40] one has $\rm(v)\Rightarrow\rm(vi)\Rightarrow(i)$. According to Remark 4.1 we know that ${\rm(ii)}\Leftrightarrow{\rm(vii)}$. One also has from [28, Theorem 5.6] that ${\rm(iv)}\Leftrightarrow{\rm(viii)}$. This completes the proof. ∎
@@
Note that the eight equivalent conditions listed in Theorem 4.1 are not exhaustive. For example, one additional equivalent condition, according to [42, Remark 3.1], is that $F$ is a locally Lipschitz homeomorphism [14, Definition 2.1.9] near $(\bar{x};\bar{y})$. Furthermore, by [42, Theorem 4.1], another condition is that $\bar{x}$ is nondegenerate and strongly stable [6, Definition 5.33]. For more details and equivalent conditions, one may refer to [42] and the references therein.
@@
According to the equivalence between the two conditions (v) and (vi) in Theorem 4.1, one can see from Lemma 4.1 that (vi) is a sufficient condition for
$0\in{\mathcal{D}}_{*}\Phi(\bar{x},\bar{y})(z)\,\Rightarrow\,z=0\quad\forall z\in{\mathcal{X}}\times{\mathcal{Y}}.$

### v2

@@ 3.2 Aubin property implies SSOSC
Since the feasible set of (3.23) is compact, the minimum of the objective function can be attained at a certain solution $(\bar{\eta};\bar{\Theta}_{1};\dots;\bar{\Theta}_{J})$. Moreover, by applying Lemma 2.5 with ${\mathcal{H}}={\mathcal{A}},\,K=-({\mathcal{S}}^{|\beta_{1}|}_{+}\times\dots\times{\mathcal{S}}^{|\beta_{J}|}_{+}),\,{\mathcal{E}}={\mathds{R}}^{r}$ and ${\mathcal{F}}={\mathcal{S}}^{|\beta_{1}|}\times\dots\times{\mathcal{S}}^{|\beta_{J}|}$, one has that Robinson’s constraint qualification holds for the constraint in (3.23) at $(\bar{\eta};\bar{\Theta}_{1};\dots;\bar{\Theta}_{J})$. Consequently, we know from [7, Theorem 3.9] that there exists a Lagrange multiplier $(\bar{\rho};\bar{\tau};\bar{\Delta}_{1};\dots;\bar{\Delta}_{J})\in{\mathds{R}}^{r}\times{\mathds{R}}\times{\mathcal{S}}^{|\beta_{1}|}\times\dots\times{\mathcal{S}}^{|\beta_{J}|}$ at $(\bar{\eta};\bar{\Theta}_{1};\dots;\bar{\Theta}_{J})$ such that the following KKT system of (3.23) holds:
$\begin{cases}({\mathcal{W}}^{*}{\mathcal{Q}}{\mathcal{W}})^{-1}\bar{\eta}+\bar{\rho}=0,\\[5.69054pt] -(P_{j})_{\beta_{j}}^{\top}(g_{j}^{\prime}(\bar{x}){\mathcal{W}}\bar{\rho})(P_{j})_{\beta_{j}}+\bar{\tau}\bar{\Theta}_{j}+\bar{\Delta}_{j}=0,\\[5.69054pt] \bar{\eta}-\sum_{j=1}^{J}{\mathcal{W}}^{*}\nabla g_{j}(\bar{x})(P_{j})_{\beta_{j}}\bar{\Theta}_{j}(P_{j})_{\beta_{j}}^{\top}=0,\\[5.69054pt] \frac{1}{2}(\sum_{j=1}^{J}\|\bar{\Theta}_{j}\|^{2}-1)=0,\\[5.69054pt] \bar{\Theta}_{j}\preceq 0,\,\bar{\Delta}_{j}\succeq 0,\langle\bar{\Delta}_{j},\bar{\Theta}_{j}\rangle=0,\;j=1,\dots,J.\end{cases}$ (3.24)
It is obvious that
@@
$\bar{\tau}$ $=\langle(\bar{\Delta}_{1};\dots;\bar{\Delta}_{J})+\bar{\tau}(\bar{\Theta}_{1};\dots;\bar{\Theta}_{J}),(\bar{\Theta}_{1};\dots;\bar{\Theta}_{J})\rangle$ (3.26)
@@
$=\Big{\langle}\bar{\rho},\sum_{j=1}^{J}{\mathcal{W}}^{*}\nabla g_{j}(\bar{x})(P_{j})_{\beta_{j}}\bar{\Theta}_{j}(P_{j})_{\beta_{j}}^{\top}\Big{\rangle}=\langle\bar{\rho},\bar{\eta}\rangle=-\langle\bar{\eta},({\mathcal{W}}^{*}{\mathcal{Q}}{\mathcal{W}})^{-1}\bar{\eta}\rangle.$
According to the last line of (3.24) one has the eigenvalue decompositions
@@
where each $\bar{O}_{j}$ is an orthogonal matrix such that $(\bar{\Lambda}_{j})_{\bar{\alpha}_{j}\bar{\alpha}_{j}}\succ 0$ and $(\bar{\Lambda}_{j})_{\bar{\gamma}_{j}\bar{\gamma}_{j}}\prec 0$. In this case, it holds that
@@
$\bar{\Delta}_{j}+\bar{\tau}\bar{\Theta}_{j}=\bar{O}_{j}\begin{pmatrix}(\bar{\Lambda}_{j})_{\bar{\alpha}_{j}\bar{\alpha}_{j}}&0&0\\ 0&0_{\bar{\beta}_{j}\bar{\beta}_{j}}&0\\ 0&0&\bar{\tau}(\bar{\Lambda}_{j})_{\bar{\gamma}_{j}\bar{\gamma}_{j}}\end{pmatrix}\bar{O}_{j}^{\top},\quad j=1,\dots,J.$ (3.27)
For an arbitrarily given index $j\in\{1,\ldots,J\}$, one can take $\beta_{+}=\bar{\alpha}_{j}$, $\beta_{0}=\bar{\beta}_{j}\cup\bar{\gamma}_{j}$ and $\beta_{-}=\emptyset$ in (2.6) and (2.7) to set
@@
$\begin{cases}\Xi_{1}\circ\bar{O}^{\top}\bar{\Theta}_{j}\bar{O}=0=\Xi_{2}\circ\bar{O}^{\top}(\bar{\Delta}_{j}+\bar{\tau}\bar{\Theta}_{j})\bar{O},\\ \bar{O}_{\beta_{0}}^{\top}\bar{\Theta}_{j}\bar{O}_{\beta_{0}}\preceq 0.\end{cases}$ (3.28)
Now, suppose on the contrary that (3.21) does not hold. With the help of the explicit formula of $\Omega^{\circ}$ in (3.22), it is easy to see that the optimal value of (3.23) is not positive, so that $\bar{\tau}\geq 0$ by (3.26). Thus, by (3.27) one has
$\bar{O}_{\beta_{0}}^{\top}(\bar{\Delta}_{j}+\bar{\tau}\bar{\Theta}_{j})\bar{O}_{\beta_{0}}\preceq 0,$
which, together with (3.28) and Lemma 2.1, implies that
$\bar{\Theta}_{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{|\beta_{j}|}}(0,0)(\bar{\Delta}_{j}+\bar{\tau}\bar{\Theta}_{j})={\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{|\beta_{j}|}}(0,0)((P_{j})_{\beta_{j}}^{\top}(g_{j}^{\prime}(\bar{x}){\mathcal{W}}\bar{\rho})(P_{j})_{\beta_{j}}),$
where the equality holds from the second line of (3.24). Note that such an inclusion holds simultaneously for all $j=1,\dots,J$. Thus, this inclusion, together with (3.25), makes a contradiction to (3.13) in Proposition 3.2 (with $\nu=\bar{\rho}$ and $\widetilde{U}^{j}_{\beta_{j}\beta_{j}}=\bar{\Theta}_{j}$ for all $j=1,\ldots,J$). Consequently, we know that (3.21) is valid, which completes the proof. ∎
4 Characterizations of the Aubin property for NLSDP
This section establishes the equivalent characterizations of the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ in (1.6) at $(0;0)$ for $(\bar{x};\bar{y})$ with $\bar{x}$ being a locally optimal solution to the NLSDP problem (1.8) and $\bar{y}\in{\mathcal{M}}(\bar{x})$.
We first review some related concepts in variational analysis. As mentioned in Section 1, the Aubin property is related to the strong metric regularity [17, Definition 2.5]. Specifically, for a set-valued mapping $\Psi:{\mathcal{E}}\rightrightarrows{\mathcal{F}}$, one has $\Psi$ is strongly metrically regular at $(\bar{z};\bar{w})\in{\sf gph}\Psi$ if $\Psi^{-1}$ has the Aubin property at $\bar{w}$ for $\bar{z}$, and there exist neighborhoods ${\mathcal{U}}$ of $\bar{z}$ and ${\mathcal{V}}$ of $\bar{w}$, such that $\Psi^{-1}(w)\cap{\mathcal{U}}$ is a singleton for all $w\in{\mathcal{V}}$. The following result provides a criterion for characterizing the strong metric regularity.
@@ 4 Characterizations of the Aubin property for NLSDP
[17, Theorem 2.7] $\Psi$ is strongly metrically regular at $(\bar{z};\bar{w})\in{\sf gph}(\Psi)$ if and only if for all $w\in{\mathcal{F}}$ and $z\in{\mathcal{E}}$, one has
@@
Since $F$ is locally Lipschitz continuous around $(\bar{x};\bar{y})$, it is almost everywhere differentiable in a neighborhood ${\mathcal{V}}$ of $(\bar{x};\bar{y})$ by Rademacher’s theorem [43, Theorem 9.60]. We use $D_{F}\subseteq{\mathcal{V}}$ to denote the set of points at which $F$ is differentiable. The Bouligand subdifferential of $F$ at $(\bar{x};\bar{y})$ is defined by
@@
Moreover, the Clarke generalized Jacobian of $F$ at $(\bar{x};\bar{y})$ is defined by
@@
$\min_{x\in{\mathcal{X}}}\ \phi(x,b)-\langle a,x\rangle,$ (4.2)
@@
$\begin{cases}\displaystyle M_{\iota}(a,b):=\operatorname*{arg\,min}_{x\in{\mathcal{X}}}\{\phi(x,b)-\langle a,x\rangle\mid\|x-\bar{x}\|\leq\iota\},\\ \displaystyle m_{\iota}(a,b):=\inf_{x\in{\mathcal{X}}}\{\phi(x,b)-\langle a,x\rangle\mid\|x-\bar{x}\|\leq\iota\}.\end{cases}$ (4.4)
We say the point $\bar{x}$ is a Lipschitzian fully stable local minimizer [30, Definition 3.2] of (4.2) at $(\bar{a};\bar{b})$ if there exist numbers $\kappa>0,\iota>0$ and a neighborhood ${\mathcal{V}}$ of $(\bar{a};\bar{b})$ such that the mapping $M_{\iota}(a,b)$ is single-valued on ${\mathcal{V}}$ with $M_{\iota}(\bar{a},\bar{b})=\bar{x}$ satisfying
@@
and that the function $m_{\iota}(a,b)$ is also Lipschitz continuous on ${\mathcal{V}}$. Unlike ${\mathbb{S}}_{\rm KKT}$, the mapping $M_{\iota}(a,b)$ focuses mainly on locally optimal solutions.
Recently, Rockafellar [41, 42] introduced the strong variational sufficient condition for local optimality and provided several characterizations of this abstract property. With $\phi$ defined in (4.3), one can reformulate (1.1) to
@@
Define the function $\phi_{\ell}(x,b):=\phi(x,b)+\frac{\ell}{2}\|b\|^{2}$. The following definition comes from [42, Section 2].
@@
and $\varrho(x,b)=\phi_{\ell}(x,b)$ holds for $((x;b);(a;y))$ belonging to this common set. Here, $\partial\varrho$ is the subdifferential in convex analysis and $\partial\phi_{\ell}$ is the limiting subdifferential defined in (2.1).
@@
For a special case ($J=1$ and $m=0$) of the NLSDP problem (1.8), the strong variational sufficient condition (with respect to a solution $(\bar{x};\bar{y})$ to the KKT system (1.3)) and the SSOSC (2.18) (for the same $(\bar{x};\bar{y})$) were proved to be equivalent in [46, Theorem 3.3]. This equivalence can be extended to the NLSDP problem (1.8) in its general form with ease by directly following their proof.
More recently, the primal-dual full stability was studied in [2] as an extension of the above (primal) full stability. Specifically, given $\bar{y}\in{\mathcal{M}}(\bar{x})$, in addition to $M_{\iota}$ in (4.4), one defines
$\overline{M}_{\iota}(a,b):=\{(x,y)\mid x\in M_{\iota}(a,b),\,(a,y)\in\partial\phi(x,b),\,\|y-\bar{y}\|\leq\iota\}.$
We say that the primal-dual pair $(\bar{x};\bar{y})$ is fully stable [2, Definition 1.4] in problem (1.8) if there exist a number $\iota>0$ and a neighborhood ${\mathcal{V}}$ of $(0;0)$ such that the mapping $\overline{M}_{\iota}$ is single-valued and Lipschitz continuous in ${\mathcal{V}}$, and the function $m_{\iota}$ is likewise Lipschitz continuous on ${\mathcal{V}}$.
Based on the above definitions, the following result holds regarding the equivalent characterizations of the Aubin property for the solution mapping ${{\mathbb{S}}}_{\rm KKT}$ in (1.6) of the NLSDP problem (1.8).
@@
    The strong second-order sufficient condition (2.18) holds at $(\bar{x};\bar{y})$ and $\bar{x}$ is nondegenerate, i.e., (2.14) holds.
@@
One has $\rm(i)\Rightarrow\rm(ii)$ from Theorem 3.1 and [25, Theorem 1]. By simply repeating the proof of [45, Proposition 3.2] one can get $\rm(ii)\Rightarrow\rm(iii)\Rightarrow\rm(iv)$. It follows by [37, Corollary 2.2] and the definition that $\rm(iv)\Rightarrow\rm(v)$ holds. From Lemma 4.1 and the Mordukhovich criterion for the Aubin property [43, Theorem 9.40] one has $\rm(v)\Rightarrow\rm(vi)\Rightarrow(i)$. According to Remark 4.1 we know that ${\rm(ii)}\Leftrightarrow{\rm(vii)}$. One also has from [30, Theorem 5.6] that ${\rm(iv)}\Leftrightarrow{\rm(viii)}$. Moreover, from [2, Theorems 2.3 & 4.2] we have $\rm(v)\Leftrightarrow\rm(ix)$. This completes the proof. ∎
In the following, we introduce an example to help with understanding Theorem 4.1, focusing on the equivalence of (i), (ii), and (iii).
Example 4.1.
Consider the following optimization problem
$\min_{x\in{\mathcal{S}}^{3}}$ $f(x):=\frac{1}{2}\|x\|^{2}-\frac{1}{2}x_{11}^{2}$ (4.5)
s.t. $h(x):=x_{11}-1=0,$
$g(x):=x\in{\mathcal{S}}_{+}^{3}.$
The unique optimal solution $\bar{x}$ of (4.5) and its unique Lagrange multiplier $\bar{y}$ are given by
$\bar{x}=\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}\quad\mbox{and}\quad\bar{y}=(\bar{\zeta},\bar{\Gamma})\quad\mbox{with}\quad\bar{\zeta}=0,\quad\bar{\Gamma}=\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}.$
It is easy to verify by definition that the constraint nondegeneracy holds at $\bar{x}$ for (4.5). By direct calculations the linear operator ${\mathcal{Q}}$ and ${\mathcal{B}}$ defined in (2.16) and (3.11) are given by
${\mathcal{Q}}d=\begin{pmatrix}0&d_{12}&d_{13}\\ d_{21}&d_{22}&d_{23}\\ d_{31}&d_{32}&d_{33}\end{pmatrix}\quad\mbox{and}\quad{\mathcal{B}}d=d_{11}\quad\forall d\in{\mathcal{S}}_{3}.$
Note that one can find a basis for ${\sf ker}{\mathcal{B}}$ given by
$\begin{array}[]{ll}\{\omega_{1},\omega_{2},\omega_{3},\omega_{4},\omega_{5}\}\\[5.69054pt] =\left\{\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{pmatrix},\begin{pmatrix}0&0&1\\ 0&0&0\\ 1&0&0\end{pmatrix},\begin{pmatrix}0&0&0\\ 0&1&0\\ 0&0&0\end{pmatrix},\begin{pmatrix}0&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix},\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&1\end{pmatrix}\right\}.\end{array}$
Taking this basis for the definition of ${\mathcal{W}}$ in (3.12), one can get
${\mathcal{W}}^{*}{\mathcal{Q}}{\mathcal{W}}\nu=(2\nu_{1};2\nu_{2};\nu_{3};2\nu_{4};\nu_{5})\quad\forall\nu=(\nu_{1};\nu_{2};\nu_{3};\nu_{4};\nu_{5})\in{\mathds{R}}^{5}.$ (4.6)
Thus the SOSC (3.20) holds, and one can use the Aubin property of ${\mathbb{S}}_{\rm GE}$ to obtain the SSOSC.
According to Proposition 3.2, the Aubin property of ${\mathbb{S}}_{\rm GE}$ at 00 for $\bar{x}$ can be equivalently expressed as that for all $0\neq\nu\in{\mathds{R}}^{5}$,
$\begin{pmatrix}2\nu_{1}\\ 2\nu_{2}\\ \nu_{3}\\ 2\nu_{4}\\ \nu_{5}\end{pmatrix}\notin\left\{\begin{pmatrix}0\\ 0\\ -u_{22}\\ -u_{23}-u_{32}\\ -u_{33}\end{pmatrix}\Big{|}\,\begin{pmatrix}u_{22}&u_{23}\\ u_{32}&u_{33}\end{pmatrix}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{2}}(0,0)\begin{pmatrix}\nu_{3}&\nu_{4}\\ \nu_{4}&\nu_{5}\end{pmatrix}\right\}.$ (4.7)
If (4.7) does not hold, there exists a certain nonzero $\nu\in{\mathds{R}}^{5}$ such that
$\nu_{1}=\nu_{2}=0\quad\mbox{and}\quad-\begin{pmatrix}\nu_{3}&\nu_{4}\\ \nu_{4}&\nu_{5}\end{pmatrix}\in{\mathcal{D}}^{*}_{{\mathcal{S}}_{+}^{2}}(0,0)\begin{pmatrix}\nu_{3}&\nu_{4}\\ \nu_{4}&\nu_{5}\end{pmatrix}.$
However, this is impossible due to [35, Theorem 2.1]. Thus, the Aubin property of ${\mathbb{S}}_{\rm GE}$ at 00 for $\bar{x}$ holds. Consequently, the analysis in Theorem 3.1 tells us that the SSOSC (3.19) holds. Of course, one can also directly observe the SSOSC from (4.6). As a result, the solution mapping ${\mathbb{S}}_{{\rm KKT}}$ also has the Aubin property at the origin for $(\bar{x};\bar{y})$.
Besides, for the function $F$ defined by (4.1), from [34, Lemma 11] and [45, Lemma 2.1] one has for any $W\in\overline{\partial}F(\bar{x},\bar{\zeta},\bar{\Gamma})$, there exists a matrix $T\in{\mathcal{S}}^{3}$ such that
$W(\Delta x,\Delta\zeta,\Delta\Gamma)=\begin{pmatrix}\begin{pmatrix}\Delta\zeta&\Delta x_{12}&\Delta x_{13}\\ \Delta x_{21}&\Delta x_{22}&\Delta x_{23}\\ \Delta x_{31}&\Delta x_{32}&\Delta x_{33}\end{pmatrix}+\Delta\Gamma\\ \Delta x_{11}\\ -\Delta x+T\circ(\Delta x+\Delta\Gamma)\end{pmatrix},$
where $\circ$” denotes the Hadamard product and
$T\in\Bigg{\{}\begin{pmatrix}1&1&1\\ 1&t_{1}&t_{2}\\ 1&t_{2}&t_{3}\end{pmatrix}\Big{|}\,t_{1},t_{2},t_{3}\in[0,1]\Bigg{\}}.$
Thus, it is easy to see that each element in ¯$\overline{\partial}F(\bar{x},\bar{\zeta},\bar{\Gamma})$ is not singular.
@@
Note that the nine equivalent conditions listed in Theorem 4.1 are not exhaustive. For example, one additional equivalent condition, according to [45, Remark 3.1], is that $F$ is a locally Lipschitz homeomorphism [15, Definition 2.1.9] near $(\bar{x};\bar{y})$. Furthermore, by [45, Theorem 4.1], another condition is that $\bar{x}$ is nondegenerate and strongly stable [7, Definition 5.33]. For more details and equivalent conditions, one may refer to [45] and the references therein.
@@
According to the equivalence between the two conditions (v) and (vi) in Theorem 4.1, one can see from Lemma 4.1 and [2, Section 2] that the conditions in Theorem 4.1 are also equivalent to
$0\in{\mathcal{D}}_{*}\Phi(\bar{x},\bar{y})(z)\quad\Rightarrow\quad z=0\quad\forall z\in{\mathcal{X}}\times{\mathcal{Y}},$
where ${\mathcal{D}}_{*}\Phi$ denotes the strict graphical derivative defined by (2.2).